# Positional Prompt Tuning for Efficient 3D Representation Learning

### v1

@@
Abstract
Point cloud analysis has achieved significant development and is well-performed in multiple downstream tasks like point cloud classification and segmentation, etc. Being conscious of the simplicity of the position encoding structure in Transformer-based architectures, we attach importance to the position encoding as a high-dimensional part and the patch encoder to offer multi-scale information. Together with the sequential Transformer, the whole module with position encoding comprehensively constructs a multi-scale feature abstraction module that considers both the local parts from the patch and the global parts from center points as position encoding. With only a few parameters, the position embedding module fits the setting of PEFT (Parameter-Efficient Fine-Tuning) tasks pretty well. Thus we unfreeze these parameters as a fine-tuning part. At the same time, we review the existing prompt and adapter tuning methods, proposing a fresh way of prompts and synthesizing them with adapters as dynamic adjustments. Our Proposed method of PEFT tasks, namely PPT, with only 1.05 of parameters for training, gets state-of-the-art results in several mainstream datasets, such as 95.01% accuracy in the ScanObjectNN OBJ_BG dataset. Codes will be released at https://github.com/zsc000722/PPT.
Introduction
@@
With the increasing popularity of scanning devices such as RGBD cameras and liDAR, we are witnessing the rise of a new modality of data: 3D point clouds. It has a wide range of applications for many tasks, such as autonomous driving and robot grasping, etc. Because of the huge potential for these downstream tasks, technological updates in 3D point cloud representation learning are iterating very fast, from PointNet (Qi et al. 2017a), PointNet++ (Qi et al. 2017b) and then all the way down to the Transformer based methods (Pang et al. 2022; Yu et al. 2022; Zhang et al. 2022; Zhao et al. 2021; Qi et al. 2023a) which now occupying the dominant position in this field. Within all these Transformer-based 3D representation learning methods, Position Encoding plays an important role in offering location specificity for Transformer architecture which can not distinguish tokens from different positions. Different from other domains like vision or language, the position encoding in the point cloud domain usually is not designed elaborately, like the rotary position embedding(Su et al. 2024) which is now widely adopted in the NLP domain to implement relative position embedding in an absolute manner. Instead, lightweight MLP with cluster centers as input serves as the position encoding module. Noticing this circumstance, we then wonder why a simple MLP as position encoding can efficiently facilitate the performance in 3D representation learning?
@@
We hold the opinion that the position encoding MLP taking center points as input, together with the patch encoder which deals with the neighbor points around the cluster centers, compose a multi-scale information extractor. Then the multi-scale patch embeddings of points are fed into the Transformer for feature subsequent feature extraction. The Max or Mean Pooling operation doesn’t break the disordered and continuous property. Thus this multi-scale method which focuses on both local and global features gains excellent performance in 3D representation learning. As shown in Figure 1(b), the average attention distance of position encoding patterns is relatively centralized while the local patch patterns hold a dispersed attention distribution, indicating the diverse concern levels of position encoding and patch tokens. In the meantime, the position encoding itself only needs a few parameters (0.2), so it is suitable for promoting the performance of PEFT (Parameter-Efficient Fine-tuning) tasks.
@@
PEFT methods usually freeze most parameters of pre-trained models, only a few selected parameters and some other ones inserted are trainable during tuning, therefore significantly releasing the demand for computational resources of full fin-tuning. In the meantime by only partly saving the trainable parameters, the pressure of storage space is also released. For language or vision models, there are Adapter tuning (Houlsby et al. 2019; Li et al. 2023), Prompt tuning (Lester, Al-Rfou, and Constant 2021; Jia et al. 2022a), and LoRA (Hu et al. 2021), which introduces extra trainable layers or prompts into pre-trained models and gets results comparable or even better than full fine-tuning. So it is natural to implement the same approach on 3D point cloud models. We attach importance to the positional embedding of point cloud representation as well as fine-tuning, unfreezing the position encoding part in the fine-tuning process, and thus get better results.
@@
In the meantime, we draw our attention to plenty of different prompt-based PEFT methods in the point cloud analysis domain (Zhou et al. 2024; Zha et al. 2023), which shares a similar paradigm. Each one of these methods inserts prompts into the encoder inputs. Instead of static prompt tokens that are selected manually, these prompts are dynamically generated by elaborate structures like the DGCNN layer in IDPT, or linear layers followed by activate function in DAPT. Both of them use an adapter-like way to generate extra prompts from the Transformer layer outputs, which brings them satisfactory results. Nevertheless, considering the prompts in the NLP domain, they are linguistically meaningful, like a series of adjectives depicting the fine-grained characteristic of the object or a specific description for the exact question. The aforementioned adapter-generated prompts are more likely to aggregate the information, thus there is nothing new for the model to get from the prompts. Nevertheless, our PPT adopts encoded sampling centers, together with the trainable positional embedding as extra prompt tokens, as shown in Figure 2. There is a fundamental distinction between this form of prompts and the previously discussed aggregate-based prompts. These prompts, obtained through sampling and clustering methods, inherently carry physical meaning in real space: specifically, the three-dimensional positional information of points. By inputting such prompts into subsequent networks and dynamically adjusting them through adapters between the Transformer layers, we can better generalize the information from pre-trained encoders to downstream tasks during training.
Our entire network architecture is remarkably straightforward and can be seamlessly integrated into transferring tasks of any Transformer-based point cloud pre-trained model. This indicates our structure is highly reusable and holds significant potential for improvement.
@@
Considering these perspectives, we propose our Positional Prompt Tuning, namely PPT. With only a few trainable parameters, we emphasize the importance of positional embedding layers and set them trainable, so as the initial encoder which transfers point clouds into patch tokens. We also insert simple trainable adapter layers between each layer of the Transformer encoder to adjust the weighted features dynamically.
@@
We revisit the Parameter-efficient Prompt Tuning problem in a Prompt and Adapter manner, and propose a quite simple method, Positional Prompt Tuning (PPT), combining the above two points.
@@
Related Works
3D Representation Learning
@@
3D Representation Learning includes point-based (Qi et al. 2017a, b), voxel-based (Maturana and Scherer 2015), and multiview-based methods (Su et al. 2015; Hamdi, Giancola, and Ghanem 2021), etc. Due to the sparse but geometry-informative representation, point-based methods (Qian et al. 2022; Engel, Belagiannis, and Dietmayer 2021) have become mainstream approaches in object classification (Wu et al. 2015; Uy et al. 2019). Voxel-based CNN methods (Yi et al. 2017; Deng et al. 2021) provide dense representation and translation invariance, achieving outstanding performance in object detection (Dai et al. 2017) and segmentation (Yi et al. 2016; Armeni et al. 2016). Furthermore, due to the vigorous development of attention mechanisms (Vaswani et al. 2017), 3D Transformers (Engel, Belagiannis, and Dietmayer 2021; Liu et al. 2021; Mao et al. 2021) have also brought about effective representations for downstream tasks. Recently, 3D self-supervised representation learning has been widely studied. PointContrast (Xie et al. 2020) leverages contrastive learning across different views to acquire discriminative 3D scene representations. Point-BERT (Yu et al. 2022) and Point-MAE (Pang et al. 2022) first introduce masked modeling pretraining into 3D. ACT (Dong et al. 2023) pioneers cross-modal geometry understanding via 2D/language foundation models. ReCon (Qi et al. 2023a, 2024) further proposes to unify generative and contrastive learning. Facilitated by foundation vision-language models like CLIP (Radford et al. 2021), another line of works are proposed towards open-world 3D representation learning (Xue et al. 2023; Peng et al. 2023; Zhang, Dong, and Ma 2023; Zhang et al. 2023a; Ding et al. 2023; Fan et al. 2023; Qi et al. 2023b).
Parameter-Efficient Fine-Tuning
@@
Pre-training and fine-tuning paradigm has become a mainstream approach in large models to maximize their semantic understanding ability for solving downstream tasks in multiple subdivisions. However, the full fine-tuning paradigm used in most cases consumes considerable storage, and may also lead to catastrophic forgetting, degrading model performance. To solve this problem, researchers in the field of NLP and 2D have proposed many effective methods as PEFT (Parameter-Efficient Fine-tuning). Prompt Tuning (Lester, Al-Rfou, and Constant 2021) and Prefix tuning (Li and Liang 2021) introduce extra inputs or vectors as the only trainable parts, while Adapter tuning (Houlsby et al. 2019) inserts additional modules between layers as the trainable part. LoRA (Hu et al. 2021) adopts low-rank approximation matrixes in a parallel manner to simulate the full fine-tuning process. VPT (Jia et al. 2022b) then introduces prompt tuning modules into image pre-trained models. These methods have achieved substantial achievements in their own specialty. So far in the 3D domain, several methods discussing the PEFT on point cloud pre-train models demonstrate their effectiveness. IDPT (Zha et al. 2023) generates instance-aware dynamic prompts instead of static prompts with a DGCNN module. DAPT (Zhou et al. 2024) then caps every layer with a light-weight TFTS layer and uses dynamic adapters to generate prompts, internally inserts in the inputs of encoders. Point-PEFT (Tang et al. 2024) constructs a point-prior bank to store feature templates of training data and uses a parameter-free attention mechanism for feature aggregation to generate prompts into first encoder layers.
PPT: Positional Prompt Tuning
@@
The existing methods of PEFT mostly summarize and transfer some existing ways from other fields to the current domain, and improve them to get better results. We attach importance to positional embedding in 3D missions and emphasize the role of position information in 3D tasks and then revisit the PEFT problem from the perspective of prompts and adapters. With several simple modifications of the network, we propose our PPT, Positional Prompt Tuning method which comprehensively considers both two aspects above, and achieves excellent results. In this section, we discuss the specifics of our PPT as follows.
Why does positional prompt matter?
@@
Positional encoding has always been a vital part of all Transformer-based structures, for their input form naturally lacks positional information. Unlike the comprehensive research in NLP (Su et al. 2024) and 2D vision (Wu et al. 2021) field, the positional encoding pattern is not as well-designed for the point cloud domain. Mostly in the 3D domain, the positional information is aggregated by passing patch centers into a simple MLP and added to the input patch tokens (Pang et al. 2022; Yu et al. 2022). Since the input of the position encoding layer is the cluster centers of the target point cloud, this highly semantic-rich input form can achieve excellent results barely with a simple MLP layer, even better than some commonly used position encoding forms in NLP, such as sinusoidal position encoding or RoPE. In fact, after re-examining the location information in 3D coped with MLP, we find that this form of position encoding is actually a multi-scale feature combined with the information of patch tokens, or to say, is a little PointNet. Due to the disorder property of point cloud, the current mainstream point cloud feature extraction network, such as PointNet (Qi et al. 2017a), generally adopts the following structure when a point cloud set is given:
@@
Where usually represents MLP or convolutional layer as point level feature extractor and denotes max or mean pooling layer. Due to the disordered and continuous characteristics of point clouds, a sequence-independent feature aggregation layer like max or mean pooling is usually adopted as adding and maximizing don’t break this property. In the point Transformer-based methods, a group of centers’ coordinates are embedded and then added to patch token inputs as positional information. With center points selected by methods like FPS and their neighbor points are given, their basic structure are as follows:
@@
where denotes the Transformer Encoder as the feature extractor, and denotes MLP encoder layer. We separately investigate the two parts of the input of the encoder which are originally added together.
@@
The branch accepts embedded center coordinates as input, which are more global messages, and together with the branch takes in more local and detailed messages from neighbor points. Integrated these two different branches together, the structure is capable of mixing information of different scales. Thus the final output of the encoder then coped with a max or mean pooling layer, is actually a multi-scale feature extractor of point cloud, holding the same effect of both utilize and . This multi-scale feature, rich in semantic information, enables the model to understand local features and at the same time to accept global location information.
@@
| PointNet (Qi et al. 2017a) | CVPR 17 | 3.5 | 73.3 | 79.2 | 68.0 | 89.2 | 90.8 |
| PointNet++ (Qi et al. 2017b) | NeurIPS 17 | 1.5 | 82.3 | 84.3 | 77.9 | 90.7 | 91.9 |
| DGCNN (Wang et al. 2019) | TOG 19 | 1.8 | 82.8 | 86.2 | 78.1 | 92.9 | - |
| PCT (Guo et al. 2021) | ICCV 21 | 2.88 | - | - | - | 93.2 | - |
| PointMLP (Ma et al. 2022) | ICLR 22 | 12.6 | - | - | 85.40.3 | 94.5 | - |
| PointNeXt (Qian et al. 2022) | NeurIPS 22 | 1.4 | - | - | 87.70.4 | 94.0 | - |
@@
| OcCo (Wang et al. 2021) | ICCV 21 | 22.1 | 84.85 | 85.54 | 78.79 | 1k | - / 92.1 |
| Point-BERT (Yu et al. 2022) | CVPR 22 | 22.1 | 87.43 | 88.12 | 83.07 | 93.2 | 93.8 |
| Point-MAE (Pang et al. 2022) | ECCV 22 | 22.1 | 90.02 | 88.29 | 85.18 | 93.8 | 94.0 |
| Point-M2AE (Zhang et al. 2022) | NeurIPS 22 | 15.3 | 91.22 | 88.81 | 86.43 | 94.0 | - |
| ACT (Dong et al. 2023) | ICLR 23 | 22.1 | 93.29 | 91.91 | 88.21 | 93.7 | 94.0 |
| VPP (Qi et al. 2023b) | NeurIPS 23 | 22.1 | 93.11 | 91.91 | 89.28 | 94.1 | 94.3 |
| I2P-MAE (Zhang et al. 2023c) | CVPR 23 | 15.3 | 94.15 | 91.57 | 90.11 | 94.1 | - |
| ReCon (Qi et al. 2023a) | ICML 23 | 44.3 | 95.18 | 93.29 | 90.63 | 94.1 | 94.3 |
| with Parameter-Efficient Supervised Finetuning | |||||||
| Point-MAE (Pang et al. 2022) (Baseline) | ECCV 22 | 22.1 | 90.02 | 88.29 | 85.18 | 93.8 | 94.0 |
| + IDPT∗ (Zha et al. 2023) | ICCV 23 | 1.7 | 92.94/92.38 | 92.60/91.33 | 88.34/88.09 | 93.64/93.02 | 93.88/93.67 |
| + DAPT∗ (Zhou et al. 2024) | CVPR 24 | 1.1 | 92.43/91.61 | 91.91/91.31 | 88.27/87.68 | 92.99/92.83 | 93.27/93.05 |
| + PPT (Ours) | - | 1.1 | 93.63/92.99 | 92.60/92.43 | 89.00/88.41 | 93.68/93.18 | 93.88/93.51 |
| ReCon (Qi et al. 2023a) (Baseline) | ICML 23 | 22.1 | 94.32 | 92.77 | 90.01 | 92.5 | 93.0 |
| + IDPT∗ (Zha et al. 2023) | ICCV 23 | 1.7 | 93.46/92.88 | 91.74/91.37 | 88.13/87.76 | 93.64/93.28 | 93.64/93.52 |
| + DAPT∗ (Zhou et al. 2024) | CVPR 24 | 1.1 | 93.63/93.12 | 92.43/91.63 | 89.31/88.77 | 93.27/92.94 | 93.07/92.79 |
| + PPT (Ours) | - | 1.1 | 95.01/94.09 | 93.28/93.23 | 89.52/88.97 | 93.76/93.41 | 93.84/93.66 |
Better way of prompt and adapter tuning?

### v2

@@
Abstract.
We rethink the role of positional encoding in 3D representation learning and fine-tuning. We argue that using positional encoding in point Transformer-based methods serves to aggregate multi-scale features of point clouds. Additionally, we explore parameter-efficient fine-tuning (PEFT) through the lens of prompts and adapters, introducing a straightforward yet effective method called PPT for point cloud analysis. PPT incorporates increased patch tokens and trainable positional encoding while keeping most pre-trained model parameters frozen. Extensive experiments validate that PPT is both effective and efficient. Our proposed method of PEFT tasks, namely PPT, with only 1.05M of parameters for training, gets state-of-the-art results in several mainstream datasets, such as 95.01% accuracy in the ScanObjectNN OBJ_BG dataset. Codes and weights will be released at https://github.com/zsc000722/PPT.
1. Introduction
@@
With the increasing popularity of scanning devices such as RGBD cameras and LiDAR, we are witnessing the rise of a new modality of data: 3D point clouds. It has a wide range of applications for many tasks, such as autonomous driving and robot grasping, etc. Because of the huge potential for these downstream tasks, technological updates in 3D point cloud representation learning are iterating very fast, from PointNet (Qi et al., 2017a), PointNet++ (Qi et al., 2017b) and then all the way down to the Transformer based methods (Pang et al., 2022; Yu et al., 2022; Zhang et al., 2022; Zhao et al., 2021; Qi et al., 2023a; Han et al., 2024) which now occupying the dominant position in this field. Within all these Transformer-based 3D representation learning methods, Position Encoding plays an important role in offering location specificity for the Transformer architecture, which can not distinguish tokens from different positions. Different from other domains like vision or language, the position encoding in the point cloud domain is usually not designed elaborately, like the rotary position embedding(Su et al., 2024), which is now widely adopted in the NLP domain to implement relative position embedding in an absolute manner. Instead, a lightweight MLP with cluster centers as input serves as the position encoding module. Noticing this circumstance, we then wonder why a simple MLP as position encoding can efficiently facilitate the performance in 3D representation learning?
@@
We hold the opinion that the position encoding MLP, taking center points as input, together with the patch encoder, which deals with the neighbor points around the cluster centers, compose a multi-scale information extractor. Then, the multi-scale patch embeddings of points are fed into the Transformer for subsequent feature extraction. The Max or Mean Pooling operation doesn’t break the disordered and continuous property. Thus, this multi-scale method, which focuses on both local and global features, gains excellent performance in 3D representation learning. As shown in fig.˜1(b), the average attention distance of position encoding patterns is relatively centralized, while the local patch patterns hold a dispersed attention distribution, indicating the diverse concern levels of position encoding and patch tokens. In the meantime, the position encoding itself only needs a few parameters (0.2), so it is suitable for promoting the performance of PEFT tasks.
@@
PEFT methods usually freeze most parameters of pre-trained models, only a few selected parameters and some other ones inserted are trainable during tuning, therefore significantly releasing the demand for computational resources of full fin-tuning. In the meantime, by only partly saving the trainable parameters, the pressure of storage space is also released. For language or vision models, there are Adapter tuning (Houlsby et al., 2019; Li et al., 2023, 2024; Zhao et al., 2024), Prompt tuning (Lester et al., 2021; Jia et al., 2022), and LoRA (Hu et al., 2021), which introduce extra trainable layers or prompts into pre-trained models and get results comparable or even better than full fine-tuning. So it is natural to implement the same approach on 3D point cloud models. We attach importance to the positional embedding of point cloud representation as well as fine-tuning, unfreezing the position encoding part in the fine-tuning process, and thus get better results.
@@
In the meantime, we draw our attention to plenty of different prompt-based PEFT methods in the point cloud analysis domain (Zhou et al., 2024; Zha et al., 2023), which share a similar paradigm. Each one of these methods inserts prompts into the encoder inputs. Instead of static prompt tokens that are selected manually, these prompts are dynamically generated by elaborate structures like the DGCNN layer in IDPT or linear layers followed by the activation function in DAPT. Both of them use an adapter-like way to generate extra prompts from the Transformer layer outputs, which brings them satisfactory results. Nevertheless, considering the prompts in the NLP domain, they are linguistically meaningful, like a series of adjectives depicting the fine-grained characteristics of the object or a specific description for the exact question. The aforementioned adapter-generated prompts are more likely to aggregate the information, thus, there is nothing new for the model to get from the prompts.
Nevertheless, our PPT adopts encoded sampling centers, together with trainable positional embedding as extra prompt tokens, as shown in fig.˜2. There is a fundamental distinction between this form of prompt and the previously discussed aggregate-based prompts. These prompts, obtained through sampling and clustering methods, inherently carry physical meaning in real space: specifically, the three-dimensional positional information of points. By inputting such prompts into subsequent networks and dynamically adjusting them through adapters between the Transformer layers, we can better generalize the information from pre-trained encoders to downstream tasks during training. Our entire network architecture is remarkably straightforward and can be seamlessly integrated into transferring tasks of any Transformer-based point cloud pre-trained model. This indicates that our structure is highly reusable and holds significant potential for improvement.
@@
Considering these perspectives, we propose our Positional Prompt Tuning, namely PPT. With only a few trainable parameters, we emphasize the importance of positional embedding layers and set them trainable, so as the initial encoder, which transfers point clouds into patch tokens. We also insert simple trainable adapter layers between each layer of the Transformer encoder to adjust the weighted features dynamically.
@@
We revisit the Parameter-efficient Prompt Tuning problem in a Prompt and Adapter manner and propose a quite simple method, Positional Prompt Tuning (PPT), combining the above two points.
@@
2. Related Works
2.1. 3D Representation Learning
@@ 2. Related Works
3D Representation Learning includes point-based (Qi et al., 2017a, b), voxel-based (Maturana and Scherer, 2015), and multiview-based methods (Su et al., 2015; Hamdi et al., 2021), etc. Due to the sparse but geometry-informative representation, point-based methods (Qian et al., 2022; Engel et al., 2021) have become mainstream approaches in object classification (Wu et al., 2015; Uy et al., 2019). Voxel-based CNN methods (Yi et al., 2017; Deng et al., 2021) provide dense representation and translation invariance, achieving outstanding performance in object detection (Dai et al., 2017) and segmentation (Yi et al., 2016; Armeni et al., 2016). Furthermore, due to the vigorous development of attention mechanisms (Vaswani et al., 2017), 3D Transformers (Engel et al., 2021; Liu et al., 2021; Mao et al., 2021) have also brought about effective representations for downstream tasks. Recently, 3D self-supervised representation learning has been widely studied(Liang et al., 2024b; Chen et al., 2024; Tang et al., 2024b; Zha et al., 2024; Zhang et al., 2024; He et al., 2024; Ma et al., 2024b; Shen et al., 2024; Luo et al., 2024; Song et al., 2025; Sun et al., 2024b; Wang et al., 2024a). PointContrast (Xie et al., 2020) leverages contrastive learning across different views to acquire discriminative 3D scene representations. Point-BERT (Yu et al., 2022) and Point-MAE (Pang et al., 2022) first introduce masked modeling pretraining into 3D. ACT (Dong et al., 2023) pioneers cross-modal geometry understanding via 2D/language foundation models. ReCon (Qi et al., 2023a, 2024) further proposes to unify generative and contrastive learning. Facilitated by foundation vision-language models like CLIP (Radford et al., 2021), another line of works are proposed towards open-world 3D representation learning (Xue et al., 2023; Peng et al., 2023; Zhang et al., 2023a, c; Ding et al., 2023; Fan et al., 2023; Qi et al., 2023b; Xia et al., 2025).
2.2. Parameter-Efficient Fine-Tuning
@@
The pre-training and fine-tuning paradigm has become a mainstream approach in large models to maximize their semantic understanding ability for solving downstream tasks in multiple subdivisions. However, the full fine-tuning paradigm used in most cases consumes considerable storage and may also lead to catastrophic forgetting, degrading model performance. To solve this problem, researchers in the field of NLP and 2D have proposed many effective methods as PEFT (Parameter-Efficient Fine-tuning). Prompt Tuning (Lester et al., 2021) and Prefix tuning (Li and Liang, 2021) introduce extra inputs or vectors as the only trainable parts, while Adapter tuning (Houlsby et al., 2019) inserts additional modules between layers as the trainable part. LoRA (Hu et al., 2021) adopts low-rank approximation matrices in a parallel manner to simulate the full fine-tuning process. VPT (Jia et al., 2022) then introduces prompt tuning modules into pre-trained image models. These methods have achieved substantial achievements in their own specialty. So far, in the 3D domain, several methods discussing the PEFT on point cloud pre-train models demonstrate their effectiveness (Zha et al., 2023; Zhou et al., 2024; Tang et al., 2024a; Liang et al., 2024a; Han et al., 2025; Fei and Deng, 2024). IDPT (Zha et al., 2023) generates instance-aware dynamic prompts instead of static prompts with a DGCNN module. DAPT (Zhou et al., 2024) then caps every layer with a light-weight TFTS layer and uses dynamic adapters to generate prompts, internally inserting them in the inputs of encoders. Point-PEFT (Tang et al., 2024a) constructs a point-prior bank to store feature templates of training data and uses a parameter-free attention mechanism for feature aggregation to generate prompts into the first encoder layers.
3. PPT: Positional Prompt Tuning
@@
The existing methods of PEFT mostly summarize and transfer some existing ways from other fields to the current domain and improve them to get better results. We attach importance to positional embedding in 3D missions and emphasize the role of position information in 3D tasks and then revisit the PEFT problem from the perspective of prompts and adapters. With several simple modifications of the network, we propose our PPT, Positional Prompt Tuning method, which comprehensively considers both aspects above and achieves excellent results. In this section, we discuss the specifics of our PPT as follows.
3.1. Why does positional prompt matter?
@@
Positional encoding has always been a vital part of all Transformer-based structures, for their input form naturally lacks positional information. Unlike the comprehensive research in NLP (Su et al., 2024) and 2D vision (Wu et al., 2021) fields, the positional encoding pattern is not as well-designed for the point cloud domain. Mostly in the 3D domain, the positional information is aggregated by passing patch centers into a simple MLP and added to the input patch tokens (Pang et al., 2022; Yu et al., 2022). Since the input of the position encoding layer is the cluster centers of the target point cloud, this highly semantic-rich input form can achieve excellent results with barely a simple MLP layer, even better than some commonly used position encoding forms in NLP, such as sinusoidal position encoding or RoPE. In fact, after re-examining the location information in 3D coped with MLP, we find that this form of position encoding is actually a multi-scale feature combined with the information of patch tokens, or to say, is a little PointNet. Due to the disorderly property of point clouds, the current mainstream point cloud feature extraction network, such as PointNet (Qi et al., 2017a), generally adopts the following structure when a point cloud set is given:
@@
Where usually represents an MLP or a convolutional layer as point point-level feature extractor and denotes a max or mean pooling layer. Due to the disordered and continuous characteristics of point clouds, a sequence-independent feature aggregation layer like max or mean pooling is usually adopted, as adding and maximizing don’t break this property. In the point Transformer-based methods, a group of centers’ coordinates are embedded and then added to the patch token inputs as positional information. With center points selected by methods like FPS and their neighbor points are given, their basic structure are as follows:
@@
where denotes the Transformer Encoder as the feature extractor, and denotes MLP encoder layer. We separately investigate the two parts of the input of the encoder, which are originally added together.
@@
The branch accepts embedded center coordinates as input, which are more global messages, and together with the branch, it takes in more local and detailed messages from neighbor points. Integrating these two different branches together, the structure is capable of mixing information of different scales. Thus, the final output of the encoder, then copied with a max or mean pooling layer, is actually a multi-scale feature extractor of point cloud, holding the same effect as both utilize and . This multi-scale feature, rich in semantic information, enables the model to understand local features and at the same time to accept global location information.
@@
| PointNet (Qi et al., 2017a) | CVPR 17 | 3.5 | 73.3 | 79.2 | 68.0 | 89.2 | 90.8 |
| PointNet++ (Qi et al., 2017b) | NeurIPS 17 | 1.5 | 82.3 | 84.3 | 77.9 | 90.7 | 91.9 |
| DGCNN (Wang et al., 2019) | TOG 19 | 1.8 | 82.8 | 86.2 | 78.1 | 92.9 | - |
| PCT (Guo et al., 2021) | ICCV 21 | 2.88 | - | - | - | 93.2 | - |
| PointMLP (Ma et al., 2022) | ICLR 22 | 12.6 | - | - | 85.40.3 | 94.5 | - |
| PointNeXt (Qian et al., 2022) | NeurIPS 22 | 1.4 | - | - | 87.70.4 | 94.0 | - |
@@
| OcCo (Wang et al., 2021) | ICCV 21 | 22.1 | 84.85 | 85.54 | 78.79 | 1k | - / 92.1 |
| Point-BERT (Yu et al., 2022) | CVPR 22 | 22.1 | 87.43 | 88.12 | 83.07 | 93.2 | 93.8 |
| Point-MAE (Pang et al., 2022) | ECCV 22 | 22.1 | 90.02 | 88.29 | 85.18 | 93.8 | 94.0 |
| Point-M2AE (Zhang et al., 2022) | NeurIPS 22 | 15.3 | 91.22 | 88.81 | 86.43 | 94.0 | - |
| ACT (Dong et al., 2023) | ICLR 23 | 22.1 | 93.29 | 91.91 | 88.21 | 93.7 | 94.0 |
| VPP (Qi et al., 2023b) | NeurIPS 23 | 22.1 | 93.11 | 91.91 | 89.28 | 94.1 | 94.3 |
| I2P-MAE (Zhang et al., 2023d) | CVPR 23 | 15.3 | 94.15 | 91.57 | 90.11 | 94.1 | - |
| ReCon (Qi et al., 2023a) | ICML 23 | 44.3 | 95.18 | 93.29 | 90.63 | 94.1 | 94.3 |
| with Parameter-Efficient Supervised Fine-tuning | |||||||
| Point-MAE (Pang et al., 2022) (Baseline) | ECCV 22 | 22.1 | 90.02 | 88.29 | 85.18 | 93.8 | 94.0 |
| + IDPT∗ (Zha et al., 2023) | ICCV 23 | 1.7 | 92.94/92.38 | 92.60/91.33 | 88.34/88.09 | 93.64/93.02 | 93.88/93.67 |
| + DAPT∗ (Zhou et al., 2024) | CVPR 24 | 1.1 | 92.43/91.61 | 91.91/91.31 | 88.27/87.68 | 92.99/92.83 | 93.27/93.05 |
| + PPT (Ours) | ACMMM 25 | 1.1 | 93.63/92.99 | 92.60/92.43 | 89.00/88.41 | 93.68/93.18 | 93.88/93.51 |
| ReCon (Qi et al., 2023a) (Baseline) | ICML 23 | 22.1 | 94.32 | 92.77 | 90.01 | 92.5 | 93.0 |
| + IDPT∗ (Zha et al., 2023) | ICCV 23 | 1.7 | 93.46/92.88 | 91.74/91.37 | 88.13/87.76 | 93.64/93.28 | 93.64/93.52 |
| + DAPT∗ (Zhou et al., 2024) | CVPR 24 | 1.1 | 93.63/93.12 | 92.43/91.63 | 89.31/88.77 | 93.27/92.94 | 93.07/92.79 |
| + PPT (Ours) | ACMMM 25 | 1.1 | 95.01/94.09 | 93.28/93.23 | 89.52/88.97 | 93.76/93.41 | 93.84/93.66 |
3.2. Better way of prompt and adapter tuning?